a canoe can travel 12 miles upstream in the same amount of time it takes to travel 30 miles downstream. if the speed of the current is 6 miles per hour, what was the speed of the canoe in still water?
Answer by mananth(16946)

let speed of canoe in still water be x
speed of current = 6mph
..
speed with current = x+6 mph
speed against current = x-6 mph
..
time = distance /rate
12/x-6 = 30/x+6 (upstream / down stream times are same )
12(x+6)=30(x-6)
12x+72=30x-180
-18x=-252
x=14 mph